Circular No. 6023

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Amplifying the announcement on IAUC 6022, T. Herbst, K. Birkle,

U. Thiele, D. Hamilton, H. Böhnhardt, A. Fiedler, K.-H. Mantel, J. L. Ortiz, G. Calamai and A. Rickicki report that the Calar Alto 2.3-μm observations were made using the MAGIC near-infrared camera of the Max-Planck-Institut für Astronomie. A flux enhancement was clearly visible on July 16.845 UT. The fireball flared to maximum brightness within a few minutes, at which time its flux surpassed that of Io. By July 16.863 the impact zone had faded significantly, although it was still visible three hours later. The fireball was fainter at 1.7 μm and not directly visible at 1.5 μm.

A report sent to the SL9 message center by P. O. Lagage, J. A. Belmonte, G. Olofsson, A. Ulla and others for the Nordic Optical Telescope (NOT) French-Swedish-Spanish team, corrects and clarifies the report on IAUC 6022: “The NOT observation of a bright spot was made at 10 μm with the Saclay CAMIRAS camera on July 16.851 UT at the southeast border of Jupiter. This spot was brighter than any pixel of Jupiter and disappeared within ∼ 30 min. A similar effect was seen at 2.16 μm with the 1.5-m Carlos Sanchez telescope at the Teide Observatory. Some two hours later, a black dot was seen in the visible video image, exactly at the position where the Sanchez telescope detected a dusty bright spot.”

J. Beish, Miami, FL, reports that observations by C. Hernandez, D. Parker and himself under very good conditions (0.40-m reflector, 381×) beginning July 16.983 UT showed a bluish cloud of appearance unprecedented in their experience near Jupiter’s south tropical belt, its core similar in size to the shadow of Ganymede.

H. B. Hammel, for the Hubble Space Telescope (HST) Team, reports that HST detected a plume on the limb of the planet starting on July 16.844 UT, and the plume had definite spatial extent above the terminator. One HST orbit later, on July 16.910, the impact site was clearly resolved, being detected in wavelengths from 336 to 953 nm and with structure. The impact site was dark at all wavelengths, except at the 889-nm methane band, where it was strikingly bright.

M. Kidger, Instituto de Astrofisica de Canarias, reports: “We can confirm the existence of a visible spot at the impact site, seen as a dark spot with a 0.82-m reflector in filters at 370, 450 and 727 nm (passbands 15, 15 and 20 nm, respectively). In the deep methane absorption band (892 nm, 5-nm passband) it is seen as a bright spot. The spot was detected here on July 16.958 UT and crossed the central meridian around July 16.979, at which time it was very bright at 892 nm.”

1994 July 16 Brian G. Marsden